OpenStudy (anonymous):

One canned juice drink is 30% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 25L that is 8% orange juice.

OpenStudy (anonymous):

x=liter of 30 percent solution y= liter of 5 percent solution .3x+.05y=25(.08) x+y=25

OpenStudy (anonymous):

How do I determine how many liters of the 30% would be used and how many of the 5% would be used?

OpenStudy (anonymous):

you would find x and y

OpenStudy (anonymous):

Then you can use substitution or set up a system of equations to solve for x and y.

OpenStudy (anonymous):

Use either x=25-y y=25-x